A $11 million project is expected to return $19 million next year. Your firm is in a 35% combined federal and state marginal income tax bracket. You finance the project with $8 million in debt at a rate of 6%. If the appropriate project interest rate is 11%, what is the present value of the tax savings due to financing the project with debt?

a. 151,351

b. 159,676

c. 319,351

d. 168,000                                                      
Answer #1
a. 151,351
Dear Student
Thank you for using Chegg
Please find below the answer
Statement showing Computations
Particulars Amount
Amount of debt    8,000,000.00
Interest Amount= 8000,000*6%        480,000.00
Tax Savings on interest = 480,000*35%        168,000.00
present value of the tax savings= 168,000/1.11        151,351.35